\begin{thebibliography}{99}
\bibitem{Fe} Fejér L., Untersuchungen uber Fouriersche Reihen, Math.
Annalen, 58 (1904), 501--569.

\bibitem{Gab} Gabisonia O. D., On strong summability points for Fourier
series, Mat. Zametki. 5, 14 (1973), 615--626.

\bibitem{GGKAM} Gát G, Goginava U., Karagulyan G., Almost everywhere strong
summability of Marcinkiewicz means of double Walsh-Fourier series. Anal.
Math. 40 (2014), no. 4, 243--266.

\bibitem{GGKJMAA} Gát G., Goginava U., Karagulyan G., On everywhere
divergence of the strong $\Phi $-means of Walsh-Fourier series. J. Math.
Anal. Appl. 421 (2015), no. 1, 206--214.

\bibitem{GGCA1} Goginava U., Gogoladze L., Strong approximation by
Marcinkiewicz means of two-dimensional Walsh-Fourier series, Constr. Approx.
35 (2012), no. 1, 1--19.

\bibitem{GGSMH} Goginava U., Gogoladze L., Strong approximation of
two-dimensional Walsh-Fourier series. Studia Sci. Math. Hungar. 49 (2012),
no. 2, 170--188.

\bibitem{GGKCA2} Goginava U., Gogoladze L., Karagulyan G., BMO-estimation
and almost everywhere exponential summability of quadratic partial sums of
double Fourier series. Constr. Approx. 40 (2014), no. 1, 105--120.

\bibitem{Gog} Gogoladze L. D., On strong summability almost everywhere.
(Russian) Mat. Sb. (N.S.) 135(177) (1988), no. 2, 158--168, 271; translation
in Math. USSR-Sb. 63 (1989), no. 1, 153--16

\bibitem{H-L} Hardy G. H., Littlewood J. E., Sur la series de Fourier d'une
fonction a carre sommable, Comptes Rendus (Paris) 156 (1913), 1307--1309.

\bibitem{JMZ} Jessen B., Marcinkiewicz J., Zygmund A., Note on the
differentiability of multiple integrals. Fund. Math. 25 (1935), 217-234.

\bibitem{Ka1}  Karagulyan G. A., On the divergence of strong $\Phi$-means of
Fourier series, Izv. Acad. Sci. of Armenia, 26(1991), no. 2, 159--162.

\bibitem{Ka} Karagulyan G. A., Everywhere divergent $\Phi $-means of Fourier
series. (Russian) Mat. Zametki 80 (2006), no. 1, 50--59; translation in
Math. Notes 80 (2006), no. 1-2, 47--56.

\bibitem{KrRu} Krasnoselski M. A., Rutitski Ya. B., Convex functions and
Orlicz spaces, Moscow, 1958 (Russian).

\bibitem{Le1} Leindler L., Über die Approximation im starken Sinne, Acta
Math. Acad. Hungar, 16 (1965), 255--262.

\bibitem{Le2} Leindler L., On the strong approximation of Fourier series,
Acta Sci. Math. (Szeged) 38 (1976), 317--324.

\bibitem{Le3} Leindler L., Strong approximation and classes of functions,
Mitteilungen Math. Seminar Giessen, 132 (1978), 29--38.

\bibitem{Le4} Leindler L., Strong approximation by Fourier series, Akademiai
Kiado, Budapest, 1985.

\bibitem{Le} Lebesgue H., Recherches sur la sommabilite forte des series de
Fourier, Math. Annalen 61 (1905), 251--280.

\bibitem{Ma2} Marcinkiewicz J., Sur une methode remarquable de sommation des
series doublefes de Fourier. \textit{Ann. Scuola Norm. Sup. Pisa}, 8(1939),
149--160.

\bibitem{Os} Oskolkov K. I., Strong summability of Fourier series. (Russian)
Studies in the theory of functions of several real variables and the
approximation of functions. Trudy Mat. Inst. Steklov. 172 (1985), 280--290.

\bibitem{Ro} Rodin, V. A., The space BMO and strong means of Fourier series.
Anal. Math. 16 (1990), no. 4, 291--302.

\bibitem{Sch} Schipp F., On the strong summability of Walsh series.
Dedicated to Professors Zoltán Daróczy and Imre Kátai. Publ. Math. Debrecen
52 (1998), no. 3-4, 611--633.

\bibitem{Sch1} Schipp F., Über die starke Summation von Walsh-Fourier
Reihen, Acta Sci. Math. (Szeged), 30 (1969), 77--87.

\bibitem{Sch2} Schipp F., On strong approximation of Walsh-Fourier series,
MTA III. Oszt. Kozl. 19(1969), 101--111 (Hungarian).

\bibitem{Sch3} Schipp F., Ky N. X., On strong summability of polynomial
expansions, Anal. Math. 12 (1986), 115--128.

\bibitem{tot} Totik V., On the strong approximation of Fourier series. Acta
Math. Acad. Sci. Hungar. 35 (1980), no. 1-2, 151--172.

\bibitem{To1} Totik V., On the strong approximation of Fourier series, Acta
Math. Sci. Hungar 35 (1980), 151--172.

\bibitem{To2} Totik V., On the generalization of Fejér's summation theorem,
Functions, Series, Operators; Coll. Math. Soc. J. Bolyai (Budapest) Hungary,
35, North Holland, Amsterdam-Oxford-New-Yourk, 1980, 1195--1199.

\bibitem{To3} Totik V., Notes on Fourier series: Strong approximation, J.
Approx. Theory, 43 (1985), 105--111.

\bibitem{WeC2} Weisz F., Strong summability of more-dimensional
Ciesielski-Fourier series. East J. Approx. 10 (2004), no. 3, 333--354.

\bibitem{WeMa} Weisz F., Lebesgue points of double Fourier series and strong
summability. J. Math. Anal. Appl. 432 (2015), no. 1, 441--462.

\bibitem{WeFAA} Weisz F., Lebesgue points of two-dimensional Fourier
transforms and strong summability. J. Fourier Anal. Appl. 21 (2015), no. 4,
885--914.

\bibitem{WeJFS} Weisz F., Strong summability of Fourier transforms at
Lebesgue points and Wiener amalgam spaces. J. Funct. Spaces 2015, Art. ID
420750, 10 pp.

\bibitem{Zy2} Zygmund A., Trigonometric series. Cambridge University Press,
Cambridge, 1959.
\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{amsmath}
\usepackage[active]{srcltx}
\usepackage{t1enc}
\usepackage[latin2]{inputenc}
\usepackage{verbatim}
\usepackage{amsmath,amsfonts,amssymb,amsthm}
\usepackage[mathcal]{eucal}
\usepackage{enumerate}
\usepackage[centertags]{amsmath}
\usepackage{graphics}

\newtheorem*{theorem}{Theorem}

\newtheorem{lemma}{Lemma}

\newtheorem*{Rodin}{Theorem R}

\newtheorem*{Kara}{Theorem K}

\def\ZT{\ensuremath{\mathbb T}}
\newcommand {\e }[1]{(\ref{#1})}

\begin{document}
\author{Ushangi Goginava and Grigori Karagulyan}
\address{U. Goginava, Department of Mathematics, Faculty of Exact and
Natural Sciences, Ivane Javakhishvili Tbilisi State University,
Chavcha\-vadze str. 1, Tbilisi 0179, Georgia}
\email{zazagoginava@gmail.com}
\address{G. A. Karagulyan, Faculty of Mathematics and Mechanics, Yerevan
State University, Alex Manoogian, 1, 0025, Yerevan, Armenia}
\email{g.karagulyan@ysu.am}
\title[Almost everywhere exponential summability]{ On almost everywhere
exponential summability of rectangular partial sums of double trigonometric
Fourier series}
\date{}
\maketitle

\begin{abstract}
In this paper we study the a.e. exponential strong summability problem for
the rectangular partial sums of double trigonometric Fourier series of the
functions from $L\log L$ .
\end{abstract}

\footnotetext{%
2010 Mathematics Subject Classification 42C10 .
\par
Key words and phrases: Double Fourier series, strong summability,
exponential means.}

\section{Introduction}

We denote the set of all non-negative integers by $\mathbb{N}$. Let $\mathbb{%
T}:=[-\pi ,\pi )=\mathbb{R}/2\pi $ and $\mathbb{R}:=\left( -\infty ,\infty
\right) $. Denote by $L^{1}\left( \mathbb{T}\right) $ the class of all
measurable functions $f$ on $\mathbb{R}$ that are $2\pi $-periodic and
satisfy 
\begin{equation*}
\left\Vert f\right\Vert _{1}:=\int\limits_{\mathbb{T}}\left\vert
f\right\vert <\infty .
\end{equation*}%
The Fourier series of a function $f\in L^{1}\left( \mathbb{T}\right) $ with
respect to the trigonometric system is 
\begin{equation}
\sum_{n=-\infty }^{\infty }c_{n}e^{inx},  \label{fourier}
\end{equation}%
where 
\begin{equation*}
c_{n}:=\frac{1}{2\pi }\int\limits_{\mathbb{T}}f\left( x\right) e^{-inx}dx
\end{equation*}%
are the Fourier coefficients of $f$. Denote by $S_{n}(x,f)$ the partial sums
of the Fourier series of $f$ and let 
\begin{equation*}
\sigma _{n}(x,f)=\frac{1}{n+1}\sum_{k=0}^{n}S_{k}(x,f)
\end{equation*}%
be the $(C,1)$ means of (\ref{fourier}). Fejér \cite{Fe} proved that $\sigma
_{n}(f)$ converges to $f$ uniformly for any $2\pi $-periodic continuous
function. Lebesgue in \cite{Le} established almost everywhere convergence of 
$(C,1)$ means if $f\in L^{1}(\mathbb{T})$. The strong summability problem,
i.e. the convergence of the strong means 
\begin{equation}
\frac{1}{n}\sum\limits_{k=0}^{n-1}\left\vert S_{k}\left( x,f\right) -f\left(
x\right) \right\vert ^{p},\quad x\in \mathbb{T},\quad p>0,  \label{Hp}
\end{equation}%
was first considered by Hardy and Littlewood in \cite{H-L}. They showed that
for any $f\in L^{r}(\mathbb{T})~\left( 1<r<\infty \right) $ the strong means
tend to $0$ a.e. as $n\rightarrow \infty $. The trigonometric Fourier series
of $f\in L^{1}(\mathbb{T})$ is said to be $\left( H,p\right) $-summable at $%
x\in \mathbb{T}$ if the values ({\ref{Hp}}) converge to $0$ as $n\rightarrow
\infty $. The $\left( H,p\right) $-summability problem in $L^{1}(\mathbb{T})$
has been investigated by Marcinkiewicz \cite{Ma2} for $p=2$, and later by
Zygmund \cite{Zy2} for the general case $1\leq p<\infty $.

Let $\Phi :[0,\infty )\rightarrow \lbrack 0,\infty )$, $\Phi \left( 0\right)
=0$, be a continuous increasing function. We say a series with the partial
sums $s_{n}$ strong $\Phi $-summable to a limit $s$ if 
\begin{equation*}
\lim_{n\rightarrow \infty }\frac{1}{n}\sum\limits_{k=0}^{n-1}\Phi
(|s_{k}-s|)=0.
\end{equation*}%
In \cite{Os} Oskolkov first considered the a.e strong $\Phi $-summability
problem of Fourier series with exponentially growing $\Phi $. Namely, he
proved a.e strong $\Phi $-summability of Fourier series if $\ln \Phi
(t)=O(t/\ln \ln t)$ as $t\rightarrow \infty $.

In \cite{Ro} Rodin proved

\begin{Rodin}[Rodin]
If a continuous function $\Phi :[0,\infty )\rightarrow \lbrack 0,\infty )$, $%
\Phi \left( 0\right) =0$, satisfies the condition 
\begin{equation*}
\limsup_{t\rightarrow +\infty }\frac{\ln \Phi \left( t\right) }{t}<\infty ,
\end{equation*}%
then for any $f\in L^{1}(\mathbb{T})$ the relation 
\begin{equation}
\lim\limits_{n\rightarrow \infty }\frac{1}{n}\sum\limits_{k=0}^{n-1}\Phi
(|S_{k}\left( x,f\right) -f\left( x\right) |)=0  \label{t3}
\end{equation}%
holds for a. e. $x\in \mathbb{T}$.
\end{Rodin}

Karagulyan \cite{Ka1, Ka} proved that the exponential growth in Rodin's
theorem is optimal. Moreover, it was proved

\begin{Kara}[Karagulyan]
If a continuous increasing function $\Phi :[0,\infty )\rightarrow \lbrack
0,\infty ),\Phi \left( 0\right) =0$, satisfies the condition%
\begin{equation*}
\limsup_{t\rightarrow +\infty }\frac{\ln \Phi \left( t\right) }{t}=\infty ,
\end{equation*}%
then there exists a function $f\in L^{1}(\mathbb{T})$, for which the relation%
\begin{equation*}
\limsup_{n\rightarrow \infty }\frac{1}{n}\sum\limits_{k=0}^{n-1}\Phi \left(
\left\vert S_{k}\left( x,f\right) \right\vert \right) =\infty 
\end{equation*}%
holds everywhere on $\mathbb{T}$.
\end{Kara}

In this paper we study the exponential summability problem for the
rectangular partial sums of double Fourier series. Let $f\in L^{1}(\mathbb{T}%
^{2})$ be a function with Fourier series 
\begin{equation}
\sum_{m,n=-\infty }^{\infty }c_{nm} e^{i(mx+ny)},  \label{DF}
\end{equation}%
where 
\begin{equation*}
c_{nm} =\frac{1}{4\pi ^{2}}\iint\limits_{\mathbb{T}%
^{2}}f(x_{1},x_{2})e^{-i(mx_{1}+nx_{2})}dx_{1}dx_{2}
\end{equation*}%
are the Fourier coefficients of the function $f$. The rectangular partial
sums of (\ref{DF}) are defined by 
\begin{equation*}
S_{MN}\left(f\right)=S_{MN}\left( x_{1},x_{2},f\right)
=\sum_{m=-M}^{M}\sum_{n=-N}^{N}c_{nm}e^{i(mx_{1}+nx_{2})}.
\end{equation*}%
We denote by $L\log L\left( \mathbb{T}^{2}\right) $ the class of measurable
functions $f$, with 
\begin{equation*}
\iint\limits_{\mathbb{T}^{2}}|f|\log^{+} |f|<\infty ,
\end{equation*}%
where $\log ^{+}u:=\mathbb{I}_{(1,\infty )}\log u$, $u>0$. For the
rectangular partial sums of two-dimensional trigonometric Fourier series
Jessen, Marcinkiewicz and Zygmund \cite{JMZ} has proved for any $f\in L\log
L\left( \mathbb{T}^{2}\right) $ that 
\begin{equation*}
\lim\limits_{n,m\rightarrow \infty }\frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum%
\limits_{j=0}^{m-1}\left( S_{ij}\left( x_{1},x_{2},f\right) -f\left(
x_{1},x_{2}\right) \right) =0
\end{equation*}%
for a. e. $\left( x_{1},x_{2}\right) \in \mathbb{T}^{2}$. They also showed
that for every non-negative function $\omega :[0,\infty )\rightarrow \lbrack
0,\infty )$ satisfying $\omega(t) \uparrow \infty$, $\omega \left( t\right)
\left( \log^+ t\right) ^{-1}\rightarrow 0$ as $t\rightarrow \infty $, there
exists a function $f$ such that $\left\vert f\right\vert \omega \left(
\left\vert f\right\vert \right) \in L^{1}\left( \mathbb{T}^{2}\right) $ and
the $\left(C,1,1\right) $ means of double Fourier series of $f$ diverge a.e..

The two dimensional a.e. strong rectangular $(H,p)$-summability, i.e. the
relation 
\begin{equation*}
\lim\limits_{n,m\rightarrow \infty }\frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum%
\limits_{j=0}^{m-1}\left\vert S_{ij}\left( x_{1},x_{2},f\right) -f\left(
x_{1},x_{2}\right) \right\vert ^{p}=0\text { a.e. }
\end{equation*}%
was proved by Gogoladze \cite{Gog} for $f\in L\log L\left( \mathbb{T}%
^{2}\right) $. These results show that in two dimensional case the optimal
class of functions for $\left( C,1,1\right) $ summability and strong
summability coincide. That is the class of functions $L\log L\left( \mathbb{T%
}^{2}\right) $.

We prove the following

\begin{theorem}
\label{exp} If  a continuous increasing function $\Phi :[0,\infty
)\rightarrow \lbrack 0,\infty )$, $\Phi \left( 0\right) =0$, satisfies the
condition 
\begin{equation}
\limsup_{t\rightarrow +\infty }\frac{\ln \Phi \left( t\right) }{\sqrt{t/\ln
\ln t}}<\infty ,  \label{t4}
\end{equation}%
then for any $f\in L\log L\left( \mathbb{T}^{2}\right) $ the relation 
\begin{equation}
\lim\limits_{n,m\rightarrow \infty }\frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum%
\limits_{j=0}^{m-1}\Phi \left( \left\vert S_{ij}\left( x_{1},x_{2},f\right)
-f(x_{1},x_{2})\right\vert \right) =0  \label{t1}
\end{equation}%
holds for a. e. $\left( x_{1},x_{2}\right) \in \mathbb{T}^{2}$.
\end{theorem}

As a corollary of this result we get the Gogoladze \cite{Gog} theorem on
a.e. $H^p$-summability of double Fourier series. From Jessen, Marcinkiewicz
and Zygmund \cite{JMZ} theorem it follows that the class $ L\log L\left( 
\mathbb{T}^{2}\right) $ in our theorem is necessary  in the context of
strong summability question. That is, it is not possible  to give a larger
convergence space than $L\log L\left( \mathbb{T} ^{2}\right) $. Our method
of proof do not allow to get (\ref{t1}) under the weaker condition  
\begin{equation}  \label{t6}
\limsup_{t\rightarrow +\infty }\frac{\ln \Phi \left( t\right) }{\sqrt t}%
<\infty.
\end{equation}
There is a conjecture that \e {t6} is the optimal bound of $\Phi$ ensuring
a.e. rectangular strong summability (\ref{t1}) for every function $f\in
L\log L\left( \mathbb{T}^{2}\right) $.

The results on strong summability and approximation by trigonometric Fourier
series have been extended for several other orthogonal systems, see Schipp 
\cite{Sch1,Sch2,Sch3}, Leindler \cite{Le1,Le2,Le3,Le4}, Totik \cite%
{tot,To1,To2,To3}, Goginava, Gogoladze \cite{GGCA1,GGSMH}, Goginava,
Gogoladze, Karagulyan \cite{GGKCA2}, Gat, Goginava, Karagulyan \cite%
{GGKAM,GGKJMAA}, Weisz \cite{WeC2}-\cite{WeJFS}.

\section{Auxiliary lemmas}

The notation $a\lesssim b$ will stand for $a<c\cdot b$, where $c>0$ is an
absolute constant. We shall write $a\sim b$ if the relations $a\lesssim b$
and $b\lesssim a$ hold at the same time. Everywhere below $q>1$ will be used
as the conjugate of $p>1$, that is $1/p+1/q=1$. $[a]$ denotes the integer
part of $a\in\mathbb{R}  $.

The maximal function of a function $f\in L^{1}\left( \mathbb{T}\right) $ is
defined by 
\begin{equation*}
Mf\left( x\right) :=\sup\limits_{I:\,x\in I\subset \mathbb{T}}\frac{1}{%
\left\vert I\right\vert }\int\limits_{I}\left\vert f\left( y\right)
\right\vert dy,
\end{equation*}%
where $I$ is an open interval. The following one dimensional operators
introduced by Gabisonia \cite{Gab} are significant tools in the
investigations of strong summability problems: 
\begin{equation*}
G_{p}^{\left( n\right) }f(x):=\left( \sum\limits_{k=1}^{\left[ n\pi \right]
}\left( \frac{n}{k}\int\limits_{\frac{k-1}{n}}^{\frac{k}{n}}\left\vert
f\left( x+t\right) \right\vert +\left\vert f\left( x-t\right) \right\vert
dt\right) ^{q}\right) ^{1/q},
\end{equation*}%
\begin{equation*}
G_{p}f(x):=\sup\limits_{n\in \mathbb{N}}G_{p}^{\left( n\right) }f(x).
\end{equation*}%
Oskolkov's following lemma plays key role in the proof of the basic lemma.

\begin{lemma}[Oskolkov, \protect\cite{Os}]
For any family of pairwise disjoint intervals $\Delta _{k}\subset \mathbb{T}$
with centers $c_{k}$ it holds the inequality 
\begin{equation}\label{a1}
\left\vert \left\{ x\in \mathbb{T}:\,\sup_{p>1}\frac{\sum_{j}\left( 
\frac{|\Delta _{j}|}{|x-c_{j}|+|\Delta _{j}|}\right) ^{q}}{p\ln \ln (p+2)}%
>\lambda \right\} \right\vert \lesssim \exp (-c\lambda ),\,\lambda >0,
\end{equation}%
where $c>0$ is an absolute constant.
\end{lemma}
One can easily check that
\begin{equation*}
\sup_{p>1}\frac{\left(\sum_{j}\left( 
	\frac{|\Delta _{j}|}{|x-c_{j}|+|\Delta _{j}|}\right)^{q}\right)^{1/q}}{p\ln \ln (p+2)}\lesssim \left\{1,\sup_{p>1}\frac{\sum_{j}\left( 
	\frac{|\Delta _{j}|}{|x-c_{j}|+|\Delta _{j}|}\right) ^{q}}{p\ln \ln (p+2)}\right\}.
\end{equation*}
Combining this with \e {a1}, we get
\begin{equation}\label{a2}
\int_{\ZT}\sup_{p>1}\frac{\left(\sum_{j}\left( 
	\frac{|\Delta _{j}|}{|x-c_{j}|+|\Delta _{j}|}\right)^{q}\right)^{1/q}}{p\ln \ln (p+2)}\lesssim 1.
\end{equation}
\begin{lemma}
\label{mainlemma}If $f\in L^{1}\left( \mathbb{T}\right) $, then%
\begin{equation}
\left\vert \left\{ x\in \mathbb{T}:\sup\limits_{p>1}\frac{G_pf(x) }{%
p\ln\ln (p+2)}> \lambda \right\} \right\vert \lesssim \left( \frac{1}{%
\lambda }\left\Vert f\right\Vert _{1}\right) ^{1/2},\quad \lambda>0.
\label{est}
\end{equation}
\end{lemma}

\begin{proof}
It is enough to prove the same estimate for the modified operators 
\begin{equation}
G_{p}^{\prime }f(x):=\sup_{n\in \mathbb{N}}\left( \sum\limits_{k=1}^{\left[
n\pi \right] }\left( \frac{n}{k}\int\limits_{\frac{k-1}{n}}^{\frac{k}{n}%
}\left\vert f\left( x+t\right) \right\vert dt\right) ^{q}\right) ^{1/q}.
\label{GanOp2}
\end{equation}%
Using the Calderon-Zygmund lemma, for the maximal function we get the
relation 
\begin{equation}
R_{\lambda }:=\left\{ x\in \mathbb{T}:Mf\left( x\right) >\sqrt{\lambda }%
\right\} =\bigcup\limits_{k=0}^{\infty }\Delta _{k},\quad \lambda >0,
\label{rep}
\end{equation}%
where $\Delta _{k}\subset \mathbb{T}$ are disjoint open intervals such that%
\begin{align}
& \sqrt{\lambda }\leq \frac{1}{\left\vert \Delta _{k}\right\vert }%
\int\limits_{\Delta _{k}}\left\vert f\left( t\right) \right\vert dt\leq 2%
\sqrt{\lambda },  \label{two-side} \\
& \left\vert R_{\lambda }\right\vert \leq \frac{1}{\sqrt{\lambda }}%
\left\Vert f\right\Vert _{1}.  \label{one-est}
\end{align}%
Denote $\delta _{k}^{n}:=\left[ \left( k-1\right) /n,k/n\right] $ and $%
\delta _{k}^{n}\left( x\right) :=x+\delta _{k}^{n}$. Separating the terms in
the sum (\ref{GanOp2}) with $k$ satisfying $\delta _{k}^{n}\left( x\right)
\subset R_{\lambda }$, we get%
\begin{align}
G_{p}^{\prime }f(x)& \leq \sup\limits_{n\in \mathbb{N}}\left(
\sum\limits_{k:\delta _{k}^{n}\left( x\right) \subset R_{\lambda }}\left( 
\frac{n}{k}\int\limits_{\frac{k-1}{n}}^{\frac{k}{n}}\left\vert f\left(
x+t\right) \right\vert dt\right) ^{q}\right) ^{1/q}  \label{gabsplit} \\
& \qquad +\sup\limits_{n\in \mathbb{N}}\left( \sum\limits_{k:\delta
_{k}^{n}\left( x\right) \not\subset R_{\lambda }}\left( \frac{n}{k}%
\int\limits_{\frac{k-1}{n}}^{\frac{k}{n}}\left\vert f\left( x+t\right)
\right\vert dt\right) ^{q}\right) ^{1/q}  \notag \\
& :=I+II.  \notag
\end{align}

From the definition of $R_{\lambda }$ in the case $\delta _{k}^{n}\left(
x\right) \not\subset R_{\lambda }$ it follows that 
\begin{equation*}
n\int\limits_{\frac{k-1}{n}}^{\frac{k}{n}}\left\vert f\left( x+t\right)
\right\vert dt\leq \sqrt{\lambda }.
\end{equation*}%
Thus we conclude 
\begin{equation}
II\leq \sqrt{\lambda }\left( \sum\limits_{k=1}^{\infty }\frac{1}{k^{q}}%
\right) ^{1/q}\lesssim \sqrt{\lambda }\left( \frac{1}{q-1}\right)
^{1/q}\lesssim p\sqrt{\lambda }.  \label{II}
\end{equation}%
Given $x\in \mathbb{T}$ set 
\begin{equation*}
k_{i}(x)=\left\{ 
\begin{array}{llr}
\min \left\{ k:\delta _{k}^{n}\left( x\right) \subset \Delta _{i}\right\}  & %
\hbox{ if }\left\{ k:\delta _{k}^{n}\left( x\right) \subset \Delta
_{i}\right\} \neq \varnothing , &  \\ 
\infty  & \hbox { if }\left\{ k:\delta _{k}^{n}\left( x\right) \subset
\Delta _{i}\right\} =\varnothing . & 
\end{array}%
\right. 
\end{equation*}%
Denote $\widetilde{R}_{\lambda }:=\bigcup\limits_{k=1}^{\infty }3\Delta _{k}$
and take an arbitrary point $x\in \mathbb{T}\backslash \widetilde{R}%
_{\lambda }$. One can easily check that if $k_{i}(x)\neq \infty $, then 
\begin{equation*}
\Delta _{i}\ni \frac{k_{i}\left( x\right) }{n}\sim \left\vert
x-c_{i}\right\vert ,
\end{equation*}%
where $c_{i}$ is the center of the interval $\Delta _{i}$. Thus for any $%
x\notin \widetilde{R}_{\lambda }$ we obtain 
\begin{align}
I& =\sup_{n\in \mathbb{N}}\left( \sum\limits_{i=1}^{\infty
}\sum\limits_{k:\delta _{k}^{n}\left( x\right) \subset \Delta _{i}}\left( 
\frac{n}{k}\int\limits_{\delta _{k}^{n}\left( x\right) }\left\vert f\left(
t\right) \right\vert dt\right) ^{q}\right) ^{1/q}  \label{I1} \\
& \leq \sup_{n\in \mathbb{N}}\left( \sum\limits_{i=1}^{\infty }\left(
\sum\limits_{k:\delta _{k}^{n}\left( x\right) \subset \Delta _{i}}\frac{n}{k}%
\int\limits_{\delta _{k}^{n}\left( x\right) }\left\vert f\left( t\right)
\right\vert dt\right) ^{q}\right) ^{1/q}  \notag \\
& \le \sup_{n\in \mathbb{N}}\left( \sum\limits_{i=1}^{\infty }\left( 
\frac{n\left\vert \Delta _{i}\right\vert }{k_{i}\left( x\right) }\frac{1}{%
\left\vert \Delta _{i}\right\vert }\int\limits_{\Delta _{i}}\left\vert
f\left( t\right) \right\vert dt\right) ^{q}\right) ^{1/q}  \notag \\
& \lesssim \sqrt{\lambda }\sup_{n}\left( \sum\limits_{i=1}^{\infty }\left( 
\frac{n\left\vert \Delta _{i}\right\vert }{k_{i}\left( x\right) }\right)
^{q}\right) ^{1/q}  \notag \\
& \lesssim \sqrt{\lambda }\left( \sum\limits_{i=1}^{\infty }\left( \frac{%
\left\vert \Delta _{i}\right\vert }{\left\vert x-c_{i}\right\vert +|\Delta
_{i}|}\right) ^{q}\right) ^{1/q},\quad x\notin \widetilde{R}_{\lambda }. 
\notag
\end{align}

Using Chebyshev's inequality, from \e {a2}, (\ref{II}) and (\ref{I1}) it follows that 
\begin{align*}
&\left\vert \left\{ x\in \mathbb{T}\backslash \widetilde{R}_{\lambda
}:\sup\limits_{p>1}\frac{G_{p}^{\prime }f(x) }{p\ln\ln (p+2)}> \lambda
\right\} \right\vert \\
&\qquad\lesssim \left\vert \left\{ x\in \mathbb{T}\backslash \widetilde{R}%
_{\lambda }:\sqrt{\lambda} \left(1+\sup\limits_{p>1}\frac{%
\left(\sum_j\left(\frac{|\Delta_j|}{|x-c_j|+|\Delta_j|}\right)^q\right)^{1/q}%
}{p\ln\ln (p+2)}\right)\geq c\lambda \right\} \right\vert  \notag \\
&\qquad\lesssim \frac{1}{\sqrt{\lambda}}\int\limits_{\mathbb{T}%
}\sup\limits_{p>1}\frac{\left(\sum_j\left(\frac{|\Delta_j|}{%
|x-c_j|+|\Delta_j|}\right)^q\right)^{1/q}}{p\ln\ln (p+2)}dx  \notag \\
&\qquad \lesssim \frac{1}{\sqrt{\lambda}},  \notag
\end{align*}
for an appropriate absolute constant $c>0$. Applying homogeneity, one can get 
\begin{equation}
\left\vert \left\{ x\in \mathbb{T}\backslash \widetilde{R}_{\lambda
}:\sup\limits_{p>1}\frac{G_{p}^{\prime }f(x) }{p\ln\ln (p+2)}> \lambda
\right\} \right\vert\lesssim \left(\frac{\|f\|_1}{\lambda}%
\right)^{1/2},\quad \lambda>0.  \label{I2}
\end{equation}

Consequently, from (\ref{one-est})-(\ref{I2}) we get%
\begin{align*}
&\left\vert \left\{ x\in \mathbb{T}:\sup\limits_{p>1}\frac{G_{p}^{\prime
}f(x) }{p\ln\ln (p+2) }>\lambda \right\} \right\vert \\
& \qquad \le \left\vert \left\{ x\in \mathbb{T}\backslash \widetilde{R}%
_{\lambda }:\sup\limits_{p>1}\frac{G_{p}^{\prime }f(x) }{p\ln\ln (p+2)}%
>\lambda \right\} \right\vert +|\tilde R_\lambda| \\
&\qquad \lesssim \left( \frac{\left\Vert f\right\Vert_1 }{\lambda }\right)
^{1/2}+ \frac{\left\Vert f\right\Vert_1 }{\sqrt \lambda }.
\end{align*}
Again using homogeneity, we obtain \e {est}.
\end{proof}

We will need the following estimations.

\begin{lemma}[Gabisonia, \protect\cite{Gab}]
\label{gabest} If $p>1$ and $f\in L^{1}(\mathbb{T}^{2})$, then 
\begin{equation}\label{a4}
\left( \frac{1}{n}\sum\limits_{j=0}^{n-1}\left\vert S_{j}(x,f)\right\vert
^{p}\right) ^{1/p}\lesssim G_{p}^{\left( n\right) }f(x).
\end{equation}
\end{lemma}

\begin{lemma}[Schipp, \protect\cite{Sch}]
\label{sch1}If $f\in L^1(\mathbb{T} ^2)$, then 
\begin{equation}\label{a5}
\left( \frac{1}{n}\sum\limits_{j=0}^{n-1}\left\vert S_{j}(x,f) \right\vert
^{p}\right) ^{1/p} \lesssim pG_2f(x) .
\end{equation}
\end{lemma}

Rodin \cite{Ro} proved the weak $(1,1)$-type estimate for the operators $%
G_{p}f(x)$ with a fixed $p>1$. From this fact, applying a standard
argument, one can derive

\begin{lemma}[Rodin, \protect\cite{Ro}]
\label{sch2}Let $f\in L\log L\left( \mathbb{T}\right) $. Then%
\begin{equation*}
\left\Vert G_2\left( f\right) \right\Vert _{1}\lesssim 1+\int\limits_{%
\mathbb{T}}\left\vert f\right\vert \log \left\vert f\right\vert .
\end{equation*}
\end{lemma}

For any function $f\in L^{1}(\mathbb{T}^{2})$ define 
\begin{eqnarray*}
G_{p,1}(x_{1},x_{2};f) &=&G_{p}f_{x_{2}}(x_{1}),\quad
G_{p,2}(x_{1},x_{2};f)=G_{p}f_{x_{1}}(x_{2}), \\
G_{p,1}^{\left( n\right) }(x_{1},x_{2};f) &=&G_{p}^{\left( n\right)
}f_{x_{2}}(x_{1}),\quad G_{p,2}^{\left( n\right)
}(x_{1},x_{2};f)=G_{p}^{\left( n\right) }f_{x_{1}}(x_{2}),
\end{eqnarray*}%
where $f_{x_{2}}(\cdot )=f(\cdot ,x_{2})$ and $f_{x_{1}}(\cdot
)=f(x_{1},\cdot )$ are considered as functions on $x_{1}$ and $x_{2}$
respectively. Similarly one dimensional partial sums of $f(x_{1},x_{2})$
with respect to each variables will be denoted by 
\begin{equation*}
S_{n,1}(x_{1},x_{2},f)=S_{n}(x_{1},f_{x_{2}}),\quad
S_{n,2}(x_{1},x_{2},f)=S_{n}(x_{2},f_{x_{1}}).
\end{equation*}

\begin{lemma}
\label{main}If $f\in L\log L(\mathbb{T}^{2})$, then 
\begin{align*}
&\left\vert \left\{ \sup\limits_{p>1}\sup\limits_{n,m\in \mathbb{N}}%
\frac{\left( \frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum\limits_{j=0}^{m-1}\left%
\vert S_{i,j}\left( x_{1},x_{2},f\right) \right\vert ^{p}\right) ^{1/p}}{%
p^{2}\ln\ln (p+2)}>\lambda \right\} \right\vert \\
&\qquad\qquad\lesssim \left( \frac{1}{\lambda }\left( 1+\iint\limits_{%
\mathbb{T}^{2}}\left\vert f\right\vert \log ^{+}\left\vert f\right\vert
\right) \right) ^{1/2},\quad \lambda>0.
\end{align*}
\end{lemma}

\begin{proof}
Using \e {a4}, \e {a5} and generalized Minkowsi's inequality, we get 
\begin{align*}
\frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum\limits_{j=0}^{m-1}\left\vert
S_{i,j}\left( x_{1},x_{2},f\right) \right\vert ^{p}& =\frac{1}{nm}%
\sum\limits_{i=0}^{n-1}\sum\limits_{j=0}^{m-1}\left\vert S_{i,1}\left(
x_{1},x_{2},S_{j,2}\left( f\right) \right) \right\vert ^{p} \\
& \leq \frac{1}{m}\sum\limits_{j=0}^{m-1}\left( G_{p,1}^{\left( n\right)
}\left( x_{1},x_{2},\left\vert S_{j,2}\left( f\right) \right\vert \right)
\right) ^{p} \\
& \leq \left( G_{p,1}^{\left( n\right) }\left( x_{1},x_{2},\left( \frac{1}{m}%
\sum\limits_{j=0}^{m-1}\left\vert S_{j,2}\left( f\right) \right\vert
^{p}\right) ^{1/p}\right) \right) ^{p} \\
& \leq \left( G_{p,1}\left( x_{1},x_{2},\left( \frac{1}{m}%
\sum\limits_{j=0}^{m-1}\left\vert S_{j,2}\left( f\right) \right\vert
^{p}\right) ^{1/p}\right) \right) ^{p} \\
& \lesssim p^{p}\left( G_{p,1}\left( x_{1},x_{2},G_{2,2}\left( f\right)
\right) \right) ^{p}.
\end{align*}%
Hence we obtain 
\begin{align*}
\Omega & =\left\{ \left( x_{1},x_{2}\right) \in \mathbb{T}%
^{2}:\sup\limits_{p>1}\sup\limits_{n,m\in \mathbb{N}}\frac{\left( \frac{1%
}{nm}\sum\limits_{i=0}^{n-1}\sum\limits_{j=0}^{m-1}\left\vert S_{i,j}\left(
x_{1},x_{2},f\right) \right\vert ^{p}\right) ^{1/p}}{p^{2}\ln \ln (p+2)}%
>\lambda \right\}  \\
& \subset \left\{ \left( x_{1},x_{2}\right) \in \mathbb{T}%
^{2}:\sup\limits_{p>1}\frac{G_{p,1}\left( x_{1},x_{2},G_{2,2}\left(
f\right) \right) }{p\ln \ln (p+2)}>\lambda \right\},
\end{align*}%
then, applying Lemma \ref{mainlemma} and \ref{sch2}, we conclude 
\begin{align*}
\left\vert \Omega \right\vert & =\int\limits_{\mathbb{T}^{2}}\mathbb{I}%
_{\Omega }\left( x_{1},x_{2}\right) dx_{1}dx_{2}=\int\limits_{\mathbb{T}%
}dx_{2}\int\limits_{\mathbb{T}}\mathbb{I}_{\Omega }\left( x_{1},x_{2}\right)
dx_{1} \\
& \lesssim \int\limits_{\mathbb{T}}\left( \frac{1}{\lambda }\int\limits_{%
\mathbb{T}}G_{2,2}\left( x_{1},x_{2},f\right) dx_{1}\right) ^{1/2}dx_{2} \\
& \lesssim \int\limits_{\mathbb{T}}\left[ \frac{1}{\lambda }\left(
1+\int\limits_{\mathbb{T}}\left\vert f\left( x_{1},x_{2}\right) \right\vert
\log ^{+}\left\vert f\left( x_{1},x_{2}\right) \right\vert dx_{1}\right) %
\right] ^{1/2}dx_{2} \\
& \lesssim \left[ \frac{1}{\lambda }\left( 1+\int\limits_{\mathbb{T}%
^{2}}\left\vert f\left( x_{1},x_{2}\right) \right\vert \log ^{+}\left\vert
f\left( x_{1},x_{2}\right) \right\vert dx_{1}dx_{2}\right) \right] ^{1/2}.
\end{align*}
Lemma is proved.
\end{proof}

\section{\protect\bigskip Proof of Theorem \protect\ref{exp}}

Let $L_{M}:=L_{M}\left(\mathbb{T} ^2\right)$ be Orlicz space of
functions on $\mathbb{T} ^2$ generated by the Young function $M(t)=t\log ^{+}t$.
It is known that $L_{M}$ is a Banach space with respect to the Luxemburg
norm 
\begin{equation*}
\left\Vert f\right\Vert _{M}:=\inf \left\{ \lambda :\lambda
>0,\int\limits_{X}M\left( \frac{\left\vert f\right\vert }{\lambda }\right)
\leq 1\right\} <\infty .
\end{equation*}%
According to a theorem from (\cite{KrRu}, Chap. 2, theorem 9.5) we have%
\begin{equation*}
0,5\left( 1+\int\limits_{\mathbb{T}^{2}}M\left( \left\vert f\right\vert
\right) \right) \leq \left\Vert f\right\Vert _{M}\leq 1+\int\limits_{\mathbb{%
\ T}^{2}}M\left( \left\vert f\right\vert \right)
\end{equation*}%
provided $\left\Vert f\right\Vert _{M}=1$. Hence from Lemma \ref{main} we
conclude 
\begin{equation}  \label{exp-est}
\left\vert \left\{ \sup\limits_{p>1}\sup\limits_{n,m\in \mathbb{N}}\frac{
\left( \frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum\limits_{j=0}^{m-1}\left\vert
S_{i,j}\left(f\right) \right\vert ^{p}\right) ^{1/p}}{ p^2\log\log (p+2)}%
>\lambda \right\} \right\vert \lesssim \left( \frac{ \left\Vert f\right\Vert
_M}{\lambda }\right) ^{1/2}.
\end{equation}%
Indeed, at first we deduce the case of $\left\Vert f\right\Vert _M=1$, then
using a homogeneity argument, we get the inequality in the general case.

\begin{proof}[Proof of Theorem]
First we shall prove that for any $f\in L\log L(\mathbb{T}^{2})$ the relation 
\begin{equation}
\lim_{n,m\rightarrow \infty }\sup_{p>1}\frac{\left( \frac{1}{nm}%
\sum\limits_{i=0}^{n-1}\sum\limits_{j=0}^{m-1}|S_{i,j}(f)-f|^{p}\right)
^{1/p}}{p^{2}\ln \ln (p+2)}=0\text{ a.e. }.  \label{t2}
\end{equation}%
Observe that \e {t2} trivially holds for the double trigonometric
polynomials. Indeed, let $T$ be a trigonometric polynomial of degree $%
(s_{1},s_{2})$. Then we have 
\begin{align*}
& S_{i,j}\left( T\right) -T=0,\quad i\geq s_{1},j\geq s_{2}, \\
& S_{i,j}\left( T\right) -T=S_{s_{1},j}\left( T\right) -T,\quad i\geq
s_{1},\,0\leq j<s_{2}, \\
& S_{i,j}\left( T\right) -T=S_{i,s_{2}}\left( T\right) -T,\quad 0\leq
i<s_{1},\,j\geq s_{2}.
\end{align*}%
Thus for integers $n> s_{1}$ and $m> s_{2}$ we have 
\begin{align*}
\frac{1}{nm}\sum_{i=0}^{n-1}& \sum_{j=0}^{m-1}|S_{i,j}(T)-T|^{p} \\
& =\frac{1}{n}\sum_{i=0}^{s_{1}-1}\frac{1}{m}%
\sum_{j=s_{2}}^{m-1}|S_{i,s_{2}}(T)-T|^{p}+\frac{1}{m}\sum_{j=0}^{s_{2}-1}%
\frac{1}{n}\sum_{i=s_{1}}^{n-1}|S_{s_{1},j}(T)-T|^{p} \\
& \qquad +\frac{1}{nm}\sum_{i=0}^{s_{1}-1}%
\sum_{j=0}^{s_{2}-1}|S_{i,j}(T)-T|^{p} \\
& \leq \frac{1}{n}\sum_{i=0}^{s_{1}-1}|S_{i,s_{2}}(T)-T|^{p}+\frac{1}{m}%
\sum_{j=0}^{s_{2}-1}|S_{s_{1},j}(T)-T|^{p} \\
& \qquad +\frac{1}{nm}\sum_{j=0}^{s_{2}-1}%
\sum_{i=0}^{s_{2}-1}|S_{i,j}(T)-T|^{p} \\
& \leq \frac{c_{1}}{n}+\frac{c_{2}}{m},
\end{align*}%
where $c_{1}$ and $c_{2}$ are constants depended on $T$. Thus (\ref{t2})
holds if $f=T$. To prove the general case it is enough to show that the set 
\begin{equation*}
G_{\lambda }=\left\{ \limsup_{n,m\rightarrow \infty }\,\sup_{p>1}\frac{%
\left( \frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum%
\limits_{j=0}^{m-1}|S_{i,j}(f)-f|^{p}\right) ^{1/p}}{p^{2}\ln \ln (p+2)}%
>\lambda \right\} 
\end{equation*}%
has measure zero for any $\lambda >0$. Since $M\left( t\right) $ satisfies
the $\Delta _{2}$-condition, the function $f$ can be approximated by a trigonometric polynomial $T$(see 
\cite{KrRu}), that is
\begin{equation*}
\left\Vert f-T\right\Vert _{M}<\varepsilon,\quad \left\Vert f-T\right\Vert _{L^1}<\varepsilon
\end{equation*}%
Since (\ref{t2}) holds for $T$, applying (\ref{exp-est}), one can obtain 
\begin{align*}
|G_{\lambda }|& =\left\vert \left\{ \limsup_{n,m\rightarrow \infty
}\,\sup_{p>1}\frac{\left( \frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum%
\limits_{j=0}^{m-1}|S_{i,j}(f-T)-(f-T)|^{p}\right) ^{1/p}}{p^{2}\ln \ln (p+2)%
}>\lambda \right\} \right\vert  \\
& \leq \left\vert \left\{ \sup_{n,m\in \mathbb{N}}\,\sup_{p>1}\frac{%
\left( \frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum%
\limits_{j=0}^{m-1}|S_{i,j}(f-T)|^{p}\right) ^{1/p}}{p^{2}\ln \ln (p+2)}%
>\lambda /2\right\} \right\vert  \\
& \qquad +\left\vert \left\{ \sup_{p>1}\frac{|f-T|}{p^{2}\ln \ln (p+2)}%
>\lambda /2\right\} \right\vert  \\
& \lesssim \left( \frac{\left\Vert f-T\right\Vert _{M}}{\lambda }\right)
^{1/2}+\frac{\|f\|_{L^1}}{\lambda} \\
& \leq \left( \frac{\varepsilon }{\lambda }\right) ^{1/2}+\frac{\varepsilon }{\lambda }.
\end{align*}%
Since $\varepsilon >0$ can be taken arbitrarily small, we conclude that $%
\left\vert G_{\lambda }\right\vert =0$ for any $\lambda >0$ and so \e {t2}
holds. To prove \e {t1} observe that 
\begin{align}
u(s)&=\exp \left( \sqrt{\frac{s}{\ln \ln (s+2)}}\right) \label{t5}\\
&\leq v(s)=\sum_{k=1}^{\infty }\left( \frac{d}{k}\sqrt{\frac{s}{\ln \ln (k+2)}}%
\right) ^{k},\,s>1,\nonumber  
\end{align}%
for some absolute constant $d$. Indeed, if $s\geq 1$, then one can check that 
\begin{equation*}
1<\sqrt{\frac{s}{\ln \ln (s+2)}}<k(s)=\left[ \sqrt{\frac{s}{\ln \ln (s+2)}}%
\right] +1<2\sqrt{\frac{s}{\ln \ln (s+2)}},
\end{equation*}%
and therefore for enough bigger $d$ we we will have
\begin{align*}
v(s)& \geq \left( \frac{d}{k(s)}\sqrt{\frac{s}{\ln \ln (k(s)+2)}}\right)
^{k(s)} \\
& >\left( \frac{d}{2}\sqrt{\frac{\ln \ln (s+2)}{\ln \ln (k(s)+2)}}\right)
^{k(s)}>e^{k(s)} \\
& \geq u(s)
\end{align*}%
and so \e {t5}. If the function $\Phi $ satisfies \e {t4}, then one
can check that 
	\begin{equation*}
	\Phi(s)\le \exp \left(\sqrt{\frac{A\cdot s}{\ln\ln (A\cdot s+2)}}\right)=u(As),\quad s>S,
	\end{equation*}
for some positive numbers $A>1,S>1$. Consider the functions 
\begin{align*}
& \varphi _{i,j}(f)=S_{i,j}(f)-f, \\
& \varphi _{i,j}^{\ast }(f)=\left\{ 
\begin{array}{lc}
\varphi _{i,j}(f) & \hbox{ if }|\varphi _{i,j}(f)|\leq S, \\ 
0 & \hbox { if }|\varphi _{i,j}(f)|>S, \\ 
& 
\end{array}%
\right.  \\
& \varphi _{i,j}^{\ast \ast }(f)=\varphi _{i,j}(f)-\varphi _{i,j}^{\ast }(f).
\end{align*}%
From \e {t5} and the definition of $\Phi$ it follows that 
\begin{align*}
\frac{1}{nm}\sum_{i=0}^{n-1}\sum_{j=0}^{m-1}\Phi (|\varphi _{i,j}(f)|)& =%
\frac{1}{nm}\sum_{i=0}^{n-1}\sum_{j=0}^{m-1}\Phi (|\varphi _{i,j}^{\ast
}(f)|)+\frac{1}{nm}\sum_{i=0}^{n-1}\sum_{j=0}^{m-1}\Phi (|\varphi
_{i,j}^{\ast \ast }(f)|) \\
& \leq \frac{1}{nm}\sum_{i=0}^{n-1}\sum_{j=0}^{m-1}\Phi (|\varphi
_{i,j}^{\ast }(f)|)+v\left(A|\varphi
_{i,j}^{\ast \ast }(f)|\right) \\
& = \frac{1}{nm}\sum_{i=0}^{n-1}\sum_{j=0}^{m-1}\Phi (|\varphi
_{i,j}^{\ast }(f)|) \\
& \qquad +\sum_{k=1}^{\infty }\frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum%
\limits_{j=0}^{m-1}\left( \frac{d}{k}\sqrt{\frac{A\cdot |\varphi
_{i,j}^{\ast \ast }(f)|}{\ln \ln (k+2)}}\right) ^{k} \\
& \leq \frac{1}{nm}\sum_{i=0}^{n-1}\sum_{j=0}^{m-1}\Phi (|\varphi
_{i,j}^{\ast }(f)|) \\
& \quad +\sum_{k=1}^{\infty }(d\sqrt{A})^{k}\left( \sup_{p>1/2}\frac{%
\left( \frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum\limits_{j=0}^{m-1}|\varphi
_{i,j}(f)|^{p}\right) ^{1/p}}{4p^{2}\ln \ln (2p+2)}\right) ^{\frac{k}{2}}.
\end{align*}%
The second term of the last expression tends to zero almost everywhere,
since according to \e {t2} we have 
\begin{align*}
\limsup_{n,m\rightarrow \infty }\sup_{p>1/2}& \frac{\left( \frac{1}{nm}%
\sum\limits_{i=0}^{n-1}\sum\limits_{j=0}^{m-1}|\varphi _{i,j}(f)|^{p}\right)
^{1/p}}{4p^{2}\ln \ln (2p+2)} \\
& \leq \lim_{n,m\rightarrow \infty }\sup_{p>1}\frac{\left( \frac{1}{nm}%
\sum\limits_{i=0}^{n-1}\sum\limits_{j=0}^{m-1}|S_{i,j}(f)-f|^{p}\right)
^{1/p}}{p^{2}\ln \ln (p+2)}=0\text{ a.e. }
\end{align*}%
Hence, to prove \e {t1} it is enough to show the same for the first term.
From \e {t2} and Chebyshev's inequality it follows that 
\begin{align*}
r_{n,m}(x_{1},x_{2})& =\frac{\#\{i,j\in \mathbb{N}:\,0\leq i<n,\,0\leq
j<m,\,\varphi _{i,j}(x_{1},x_{2})>\varepsilon \}}{nm} \\
& \leq \frac{1}{\varepsilon }\cdot \frac{1}{nm}\sum\limits_{i=0}^{n-1}\sum%
\limits_{j=0}^{m-1}|\varphi _{i,j}(x_{1},x_{2},f)|\rightarrow 0\text{ a.e.},
\end{align*}%
where $\#C$ denotes the cardinality of a finite set $C$. Thus for a.e. $%
(x_{1},x_{2})\in \mathbb{T}^{2}$ we get 
\begin{align*}
\limsup_{n,m\rightarrow \infty }\frac{1}{nm}&
\sum_{i=0}^{n-1}\sum_{j=0}^{m-1}\Phi (|\varphi _{i,j}^{\ast
}(x_{1},x_{2},f)|) \\
& \leq \limsup_{n,m\rightarrow \infty }(r_{n,m}(x_{1},x_{2})\Phi
(S)+(1-r_{n,m}(x_{1},x_{2}))\Phi (\varepsilon )) \\
& =\Phi (\varepsilon )\text{ a.e.}.
\end{align*}%
Since $\varepsilon >0$ can be taken arbitrary small we get 
\begin{equation*}
\lim_{n,m\rightarrow \infty }\frac{1}{nm}\sum_{i=0}^{n-1}\sum_{j=0}^{m-1}%
\Phi (|\varphi _{i,j}^{\ast }(x_{1},x_{2},f)|)=0\text{ a.e.}
\end{equation*}%
and so \e {t1}.
\end{proof}

\end{document}